Simplification theorems (Boolean)

1. X+XY=X

2.X(X+Y)=X

3.(X+Y')Y=XY


1
Expert's answer
2021-12-07T20:15:35-0500

(1)

X+XY =X

LHS

X+XY =X(1+Y)

Where 1+Y=1


"X+XY=X(1+Y)=X\times1=X"

"X+XY=X"

LHS=RHS

(2)

X(X+Y)=X

LHS

"X(X+Y)=X\times X+X\times Y"

XX=X

"X(X+Y)=X+XY"

"X(X+Y)=X(1+Y)"

Where 1+Y=1

"X(X+Y)=X\times1"

X(X+Y)=X

LHS=RHS

(3)

(X+Y')Y=XY

LHS

(X+Y')Y=XY+YY'

Where YY'=0

(X+Y')Y=XY+0

(X+Y')Y=XY

LHS =RHS